# Universal projection theorems with applications to multifractal analysis and the dimension of every ergodic measure on self-conformal sets simultaneously

### v2

@@ 3.3. Further examples
3.4. Applications to multifractal analysis
Finally, let us present here an application of Theorem 3.11 in the theory of multifractal analysis. Let =${\mathcal{F}}=\{f_{i}(x)=\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be a self-similar IFS on the line, let $\underline{p}=(p_{i})_{i\in{\mathcal{A}}}$ be a probability vector and let $\mu=\underline{p}^{\mathbb{N}}$ be the corresponding Bernoulli measure on $\Sigma$. In this case, the quantity $s_{0}$ defined in (3.5) is given by the equation $\sum_{i\in{\mathcal{A}}}|\lambda_{i}|^{s_{0}}=1$ and called the similarity dimension.
Arbeiter and Patschke [AP96] studied the multifractal spectrum of self-similar measures under the open set condition. For every $q\in\mathbb{R}$, let $T(q)$ be the unique number such that
$\sum_{i\in{\mathcal{A}}}p_{i}^{q}|\lambda_{i}|^{T(q)}=1.$
Arbeiter and Patschke [AP96] showed that if the OSC holds and $(p_{i})_{i\in{\mathcal{A}}}\neq(|\lambda_{i}|^{s_{0}})_{i\in{\mathcal{A}}}$, then for every $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$
(3.11) $\dim_{H}\left\{x:d(\Pi\mu,x)=\alpha\right\}=\inf_{q\in\mathbb{R}}(\alpha q+T(q))=:T^{*}(\alpha).$
In particular $T^{*}$ is an analytic concave map.
The multifractal analysis becomes significantly more difficult, if there are overlaps between the cylinder sets. Barral and Feng [BF13, Theorem 1.3, Theorem 6.4] showed that under the transversality condition, the multifractal formalim (3.11) holds on partial interval of the domain. Let us state here their result on the line under the assumption that $s_{0}<1$. That is, under these assumptions, for Lebesgue almost every $(t_{i})_{i\in{\mathcal{A}}}$ the multifractal formalism (3.11) holds for every $\alpha\in\left[\frac{\sum_{i}p_{i}\log p_{i}}{\sum_{i}p_{i}\log|\lambda_{i}|},\frac{\sum_{i}|\lambda_{i}|^{s_{0}}\log p_{i}}{\sum_{i}|\lambda_{i}|^{s_{0}}\log|\lambda_{i}|}\right]$. Later in [BF21, Theorem 1.2, Remark 7.3], they extended this result significantly. Namely, if ${\mathcal{F}}$ satisfies the exponential separation condition and $s_{0}<1$ then for every $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\frac{\sum_{i}|\lambda_{i}|^{s_{0}}\log p_{i}}{\sum_{i}|\lambda_{i}|^{s_{0}}\log|\lambda_{i}|}\right]$ the multifractal formalism (3.11) holds. Obtaining the multifractal formalism on the full interval $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ under exponential separation remains an open problem. Using Theorem 2.7 we can show that the multifractal formalism holds on the whole domain for almost every translation parameter.
@@ 3.4. Applications to multifractal analysis
Let ${\mathcal{A}}$ be a finite set and for each $i\in{\mathcal{A}}$ fix $\lambda_{i}\in(-1,1)\setminus\{0\}$. For $t=(t_{i})_{i\in{\mathcal{A}}}\in\mathbb{R}^{{\mathcal{A}}}$ let ${\mathcal{F}}_{t}=\{f_{i}(x)=\lambda_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be a self-similar IFS on the line. Let $s_{0}=s({\mathcal{F}}_{t})$ be the similarity dimension of ${\mathcal{F}}_{t}$ and assume that $s_{0}<1$. Then the following holds for Lebesgue almost every $(t_{i})_{i\in{\mathcal{A}}}\in\mathbb{R}^{\mathcal{A}}$. For every probability vector $\underline{p}=(p_{i})_{i\in{\mathcal{A}}}$ such that $(p_{i})_{i\in{\mathcal{A}}}\neq(|\lambda_{i}|^{s_{0}})_{i\in{\mathcal{A}}}$, the multifractal formalism (3.11) holds for the self-similar measure $\Pi_{{\mathcal{F}}_{t}}(\underline{p}^{\mathbb{N}})$ and every $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log|\lambda_{i}|}\right]$ simultaneously.
See Section 8 for the proof and Theorem 8.3 for a multidimensional version.
@@ 4. Preliminaries
Let $(X,\rho)$ be a metric space. The lower and upper local dimensions of a finite Borel measure $\mu\in{\mathcal{M}}_{\mathrm{fin}}(X)$ at a point $x\in{\rm supp}\mu$ are defined as
@@ 5. Relative dimension separability
$-2\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq 2\varepsilon.$
@@
$-3\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq 3\varepsilon.$
@@ 6.1. Preliminaries on relative dimension
holds for every $x\in G_{\lambda_{0},q,\varepsilon,R}(\nu)$ and $20<r<R/$ 2. Moreover, there exists $M=M(R,\varepsilon)$ such that
@@
Containment (6.1) follows directly from the definitions of $A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)$ and $G_{\lambda_{0},q,\varepsilon,M}(\nu)$. For (6.2), fix $x\in G_{\lambda_{0},q,\varepsilon,M}(\nu)$ and $20<r<R/$ 2. If $B_{\lambda_{0}}(x,r)\cap A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)=\emptyset$, then (6.2) holds trivially. Otherwise, choose $y\in B_{\lambda_{0}}(x,r)\cap A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)$. Then
@@
hold. Similarly, if $\underline{\dim}_{H}(\mu,\rho_{\lambda_{0}})\geq q$, then for $\mu$-a.e. $x\in X$ there exists $R(x)$ such that
@@
[…]
@@
(6.17) if $\text{if }\ \rho_{\lambda_{0}}(x,y)\leq D_{0}2^{-n}\text{, then }\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq D2^{-\frac{n}{1+\xi}}.$
@@
[…]
@@
(6.21) $\begin{split}\sum\limits_{x^{\prime}\in F_{n+m}}&\sum\limits_{y^{\prime}\in F_{n+m}}\nu\left(B_{\lambda_{0}}\left(x^{\prime},\frac{1}{8}D_{0}2^{-{(n+m)}}\right)\right)\nu\left(B_{\lambda_{0}}\left(y^{\prime},\frac{1}{8}D_{0}2^{-{(n+m)}}\right)\right)\\ &\qquad\times\mathds{1}_{\left\{\frac{1}{2}D_{0}2^{-(n+1)}<\rho_{\lambda_{0}}(x^{\prime},y^{\prime})\leq 2D_{0}2^{-n}\right\}}\\ &\lesssim 2^{6\varepsilon(n+m)}\sum\limits_{x^{\prime}\in F_{n+m}}\sum\limits_{y^{\prime}\in F_{n+m}}\nu\left(B_{\lambda_{0}}\left(x^{\prime},\frac{1}{16}D_{0}2^{-{(n+m)}}\right)\right)\nu\left(B_{\lambda_{0}}\left(y^{\prime},\frac{1}{16}D_{0}2^{-{(n+m)}}\right)\right)\\ &\qquad\times\mathds{1}_{\left\{\frac{1}{2}D_{0}2^{-(n+1)}<\rho_{\lambda_{0}}(x^{\prime},y^{\prime})\leq 2D_{0}2^{-n}\right\}}.\\ \end{split}$
@@
$\begin{split}\sum\limits_{x^{\prime}\in F_{n+m}}&\sum\limits_{y^{\prime}\in F_{n+m}}\nu\left(B_{\lambda_{0}}\left(x^{\prime},\frac{1}{8}D_{0}2^{-{(n+m)}}\right)\right)\nu\left(B_{\lambda_{0}}\left(y^{\prime},\frac{1}{8}D_{0}2^{-{(n+m)}}\right)\right)\\ &\qquad\times\mathds{1}_{\left\{\frac{1}{2}D_{0}2^{-(n+1)}<\rho_{\lambda_{0}}(x^{\prime},y^{\prime})\leq 2D_{0}2^{-n}\right\}}\\ &\lesssim\nu(X)2^{-n(q-2\varepsilon)+6\varepsilon(n+m)}.\\ \end{split}$
@@
Fix $L>0$ and set $U_{L}=\{\lambda\in U:\operatorname{Lip}(\Pi_{\lambda},\rho_{\lambda})\leq L\}$. For every $\lambda_{0}\in U_{L},\ \theta>0$ there exists an open neighbourhood $U^{\prime}$ of $\lambda_{0}\in U_{L}$ and a constant $D=D(\lambda_{0},L,\theta)$ such that for every $x\ \in X,0<r\leq\delta\leq 1$ inequality
@@
[…]
@@
(6.30) $\begin{split}\eta&\left(\left\{\lambda\in U^{\prime}:\underset{y\in X}{\exists}\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq r,\ \rho_{\lambda_{0}}(x,y)\geq\delta\right\}\right)\\ &\leq\sum\limits_{i=0}^{\infty}\eta\left(\left\{\lambda\in U^{\prime}:\underset{y\in X}{\exists}\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq r,\ 2^{i}\delta\leq\rho_{\lambda_{0}}(x,y)<2^{i+1}\delta\right\}\right)\\ &\leq\sum\limits_{i=0}^{\infty}\sum\limits_{m=1}^{N_{i}}\eta\left(\left\{\lambda\in U^{\prime}:\underset{y\in B_{\lambda_{0}}(y_{i,m},r)}{\exists}\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq r\right\}\right)\\ &\leq\sum\limits_{i=0}^{\infty}\sum\limits_{m=1}^{N_{i}}\eta\left(\left\{\lambda\in U^{\prime}:\underset{y\in B_{\lambda}(y_{i,m},Hr^{\frac{1}{1+\xi}})}{\exists}\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq r\right\}\right)\\ \end{split}$
@@
[…]
@@ 7.1. Relative dimension separability in symbolic spaces
whenever the limit exists (we use the standard convention $00\log\frac{0}{0}=0\log 0=$ 0).
@@
Let $\mu\in\mathcal{M}_{\sigma}(\Sigma)$ The following hold:
@@
(7.4) $[\omega|_{n+1}]\subset B(\omega,r)\subset[\omega|_{n}]$
@@
If $\tau\in B(\omega,r)$, then $\rho(\omega,\tau)=\psi(\omega\wedge\tau)<r\leq\psi(\omega|_{n})$, and hence, by ii, $\omega|_{n}$ is a prefix of $\omega\wedge\tau$. This implies $\tau\in[\omega|_{n}]$, so $B(\omega,r)\subset[\omega|_{n}]$. On the other hand, if $\tau\in[\omega|_{n+1}]$, then $\omega|_{n+1}$ is a prefix of $\omega\wedge\tau$, so ii gives $\rho(\omega,\tau)=\psi(\omega\wedge\tau)\leq\psi(\omega|_{n+1})<r$. Therefore $[\omega|_{n+1}]\subset B(\omega,r)$. This proves (7.4). Set $A=\max\{\psi(i):i\in{\mathcal{A}}\}$ and note that ii implies
@@
$-A\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\mu(B(\omega,r))}{\nu(B(\omega,r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\mu(B(\omega,r))}{\nu(B(\omega,r))}}{\log r}\leq A\varepsilon.$

### v3

@@ 3.3. Further examples
3.4. Applications to the multifractal analysis
Finally, let us discuss in more details the application of Theorem 3.12 to the multifractal analysis of self-similar measures. Let ${\mathcal{F}}_{t}=\{f_{i}(x)=\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ be an IFS consisting of similarities on $\mathbb{R}^{d}$, so that $\lambda_{i}\in(0,1)$, $O_{i}$ is a $d\times d$ orthogonal matrix and $t_{i}\in\mathbb{R}^{d}$. Let $p=(p_{i})_{i\in{\mathcal{A}}}$ be a probability vector. For every $q\in\mathbb{R}$ we define $T(q)$ to be the unique real number satisfying
$\sum_{i\in{\mathcal{A}}}p_{i}^{q}\lambda_{i}^{T(q)}=1.$
The general version of our main result on the multifractal spectra of self-similar measures is the following.
@@ 3.4. Applications to the multifractal analysis
Fix $d\geq 1$. Let $\mathcal{A}$} be a finite set and for each $i\in{\mathcal{A}}$ fix $\lambda_{i}\in(0,1)$ and a $d\times d$ orthogonal matrix $O_{i}$. For each $t=(t_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{{\mathcal{A}}}$ and a probability vector $p=(p_{i})_{i\in{\mathcal{A}}}$, let $\nu_{t,p}$ be the self-similar measure corresponding to the IFS ${\mathcal{F}}_{t}=\{f_{i}(x)=\lambda_{i}O_{i}x+t_{i}\}_{i\in{\mathcal{A}}}$ and the probability vector $p$. The following holds for Lebesgue almost every $(t_{i})_{i\in{\mathcal{A}}}\in(\mathbb{R}^{d})^{\mathcal{A}}$. If the similarity dimension $s_{0}=s({\mathcal{F}}_{t})$ of ${\mathcal{F}}_{t}$ satisfies
  1. (i)
    $s_{0}<1$ if $d=1$, or
  2. (ii)
    $s_{0}<d$ and $\max\limits_{i\in{\mathcal{A}}}\lambda_{i}<1/2$ if $d\geq 2$,
then equality
$\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)=\inf_{q\in\mathbb{R}}(\alpha q+T(q))$
holds for
  1. (i)
    every $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$ if $d=1$, or
  2. (ii)
    every $\alpha\in\left[\frac{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log p_{i}}{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$ if $d\geq 2$.
and every probability vector $p$ such that $p_{i}\neq\lambda_{i}^{s_{0}}$ for some $i\in{\mathcal{A}}$.
We shall make now several comments on the above result.
Remark 3.16.
Theorem 3.15 with $d>1$ does not cover full range $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$. This is because our technique deals only with decreasing part of the spectrum $\left[\frac{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log p_{i}}{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log\lambda_{i}},\max\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}}\right]$, while the increasing part $\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\frac{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log p_{i}}{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log\lambda_{i}}\right]$ follows from the result of Barral and Feng [BF21, Theorem 1.2, Remark 7.3]. As the latter was so far obtained only for $d=1$, the full range result is limited to this case. Nevertheless, it follows from the proof of Theorem 3.15 that in the remaining part of the spectrum we obtain a lower bound on the dimension of the level set, i.e.
$\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)\geq\inf_{q\in\mathbb{R}}(\alpha q+T(q))$
holds for $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\frac{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log p_{i}}{\sum\limits_{i\in{\mathcal{A}}}\lambda_{i}^{s_{0}}\log\lambda_{i}}\right]$.
Remark 3.17.
We shall emphasize that Theorem 3.15, even for $d=1$, does not establish the full multifractal formalism for the corresponding self-similar measures. One says that the multifractal formalism holds at $\alpha$ if
$\dim_{H}\left(\left\{x:d(\nu_{t,p},x)=\alpha\right\}\right)=\inf_{q\in\mathbb{R}}(\alpha q-\tau_{\nu_{t,p}}(q)),$
where $q\mapsto\tau_{\nu_{t,p}}(q)$ is the $L^{q}$-spectrum of the self-similar measure $\nu_{t,p}$, see e.g. [BSS23, Chapter 5]. Our result shows that the dimension of the level sets of the local dimension is preserved under the projection from the symbolic space onto the self-similar set. Establishing the multifractal formalism in the setting of Theorem 3.15, requires proving additionally that $\tau_{\nu_{t,p}}(q)=-T(q)$ for $q\in\mathbb{R}$. For $q>0$ this was proved by Shmerkin for self-similar measures satisfying the exponential separation condition [Shm19]. The case $q<0$ remains an open problem. The results of Barral and Feng [BF21] establish the full multifractal formalism for $d=1$ in the restricted range $\alpha\in\left[\min\limits_{i\in{\mathcal{A}}}\frac{\log p_{i}}{\log\lambda_{i}},\frac{\sum\limits_{i\in{\mathcal{A}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{i}^{s_{0}}}\log p_{i}}{\sum\limits_{i\in{\mathcal{A}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{i}^{s_{0}}}\log\lambda_{i}}\right]$ (corresponding to $q>0$), for self-similar measures on $\mathbb{R}$ satisfying the exponential separation condition.
@@ 4. Preliminaries
Let $(X,\rho)$ be a metric space. The lower and upper local dimensions of a finite Borel measure $\mu\in{\mathcal{M}}_{\mathrm{fin}}(X)$ at a point $x\in{\rm supp}(\mu)$ are defined as
@@ 5. Relative dimension separability
$2\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq 2\varepsilon.$
@@
$3\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\Pi\mu(B(\Pi(x),r))}{\Pi\nu(B(\Pi(x),r))}}{\log r}\leq 3\varepsilon.$
@@ 6.1. Preliminaries on relative dimension
holds for every $x\in G_{\lambda_{0},q,\varepsilon,R}(\nu)$ and $0<r<R/2$. Moreover, there exists $M=M(R,\varepsilon)$ such that
@@
Containment (6.1) follows directly from the definitions of $A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)$ and $G_{\lambda_{0},q,\varepsilon,M}(\nu)$. For (6.2), fix $x\in G_{\lambda_{0},q,\varepsilon,M}(\nu)$ and $0<r<R/2$. If $B_{\lambda_{0}}(x,r)\cap A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)=\emptyset$, then (6.2) holds trivially. Otherwise, choose $y\in B_{\lambda_{0}}(x,r)\cap A_{\lambda_{0},q,\varepsilon,R}(\mu,\nu)$. Then
@@
hold.
Similarly, if $\underline{\dim}_{H}(\mu,\rho_{\lambda_{0}})\geq q$, then for $\mu$-a.e. $x\in X$ there exists $R(x)$ such that
@@
(6.16) $\begin{split}{\mathcal{E}}_{s}&(\Pi_{\lambda}\tilde{\mu})=\int\limits_{G}\int\limits_{G}|\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|^{-s}d\tilde{\mu}(x)d\tilde{\mu}(y)\\ &=\sum\limits_{n=0}^{\infty}\int\limits_{G}\int\limits_{G}\mathds{1}_{\left\{D_{0}2^{-(n+1)}<\rho_{\lambda_{0}}(x,y)\leq D_{0}2^{-n}\right\}}|\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|^{-s}d\tilde{\mu}(x)d\tilde{\mu}(y).\end{split}$
@@
(6.17) $\text{if }\ \rho_{\lambda_{0}}(x,y)\leq D_{0}2^{-n}\text{, then }\ |\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq D2^{-\frac{n}{1+\xi}}.$
@@
Indeed, if $\rho_{\lambda_{0}}(x,y)\leq D_{0}2^{-n}$ and $\Pi_{\lambda}(x)\neq\Pi_{\lambda}(y)$, then by (6.17) there exists $m\geq 0$ such that $2^{-(m+1)}D2^{-\frac{n}{1+\xi}}<|\Pi_{\lambda}(x)-\Pi_{\lambda}(y)|\leq 2^{-m}D2^{-\frac{n}{1+\xi}}$, so (6.18) follows. If $\Pi_{\lambda}(x)=\Pi_{\lambda}(y)$ then both sides of the inequality in (6.18) are infinite provided $D_{0}2^{-(n+1)}<\rho_{\lambda_{0}}(x,y)\leq D_{0}2^{-n}$. Applying (6.18) to (6.16) gives
@@
[…]
@@
[…]
@@
[…]
@@
It exists due to the definition of the Assouad dimension (note that $C_{\xi}$ does not depend on $x$). By (6.26) we have for $\lambda\in U^{\prime},\ i\geq 0,\ 1\leq m\leq N_{i}$
@@
[…]
@@
 $\text{ for }\eta\text{-a.e. }\lambda\in U^{\prime},\ \lim\limits_{C\to\infty}\mu|_{A_{\lambda_{0},0,\varepsilon,R}(\mu,\nu)}\left(E_{C,\lambda}\right)=0\text{ for every }\mu\in\underline{V}_{\lambda_{0},0,\varepsilon}(\nu).$
@@ 7.1. Relative dimension separability in symbolic spaces
whenever the limit exists (we use the standard convention $0\log\frac{0}{0}=0\log 0=0$).
@@
Let $\mu\in\mathcal{M}_{\sigma}(\Sigma)$ . The following hold:
@@
(7.4) $\omega|_{n+1}]\subset B(\omega,r)\subset[\omega|_{n}]$
@@
If $\tau\in B(\omega,r)$, then $\rho(\omega,\tau)=\psi(\omega\wedge\tau)<r\leq\psi(\omega|_{n})$, and hence, by ii, $\omega|_{n}$ is a prefix of $\omega\wedge\tau$. This implies $\tau\in[\omega|_{n}]$, so $B(\omega,r)\subset[\omega|_{n}]$. On the other hand, if $\tau\in[\omega|_{n+1}]$, then $\omega|_{n+1}$ is a prefix of $\omega\wedge\tau$, so ii gives $\rho(\omega,\tau)=\psi(\omega\wedge\tau)\leq\psi(\omega|_{n+1})<r$. Therefore $\omega|_{n+1}]\subset B(\omega,r)$. This proves (7.4). Set $A=\max\{\psi(i):i\in{\mathcal{A}}\}$ and note that ii implies
@@
$A\varepsilon\leq\liminf\limits_{r\to 0}\frac{\log\frac{\mu(B(\omega,r))}{\nu(B(\omega,r))}}{\log r}\leq\limsup\limits_{r\to 0}\frac{\log\frac{\mu(B(\omega,r))}{\nu(B(\omega,r))}}{\log r}\leq A\varepsilon.$